# Quake: Adaptive Indexing for Vector Search

### v1

@@ 4.2.4 Example Maintenance Workflow
    Estimate.  For $P_{1}$ and $P_{2}$ the estimate assumes a balanced $/250$ 250 split and $\alpha=0.5$ traffic per child:
@@
    Verify.  After splitting we see that $P_{1}$ has a $/250$ 250 split, but $P_{2}$ comes out $450/50$:
@@ 5 Adaptive Partition Scanning (APS)
Adaptive Partition Scanning (APS) dynamically determines the number of partitions to scan per query to achieve a specified recall target $\tau_{R}$ with minimal latency. APS adapts to evolving workloads and changing index structures, making it particularly effective in dynamic data settings. We first introduce the geometric model underlying APS, followed by a detailed description of the scanning algorithm, and conclude with key performance optimizations. We apply APS at each level of the index independently.
@@
Exact computation of intersection volumes between a sphere and high-dimensional Voronoi partition boundaries is infeasible, as partitions are intersections of multiple half-spaces. Instead, we approximate each partition as a single half-space, defined by the perpendicular bisector between the query’s nearest centroid $\mathbf{c}_{0}$ and each neighboring centroid $\mathbf{c}_{i}$. This simplification results in a hyperspherical cap whose volume $v_{i}$ has a closed-form expression via the regularized incomplete beta function [19] (see technical report).
@@
$p_{0}=\prod_{j=1}^{M-1}(1-v_{j}),$
with the remaining probability distributed proportionally among other partitions:
$p_{i}=(1-p_{0})\,v_{i}.$
Thus, these probabilities sum to 1 and provide valid recall estimates.
@@ 5.1 APS algorithm
Algorithm 1 details the APS procedure. Given query $\mathbf{q}$, recall target $\tau_{R}$, and an initial set of $M$ candidate centroids:
@@
This process is conducted at each level of the index. To avoid propagating errors from searching higher levels of the index, we fix the recall target to .999 for the higher levels.
@@
1:query $\mathbf{q}$, centroids $\mathcal{C}$, recall target $\tau_{R}$, initial candidate count $M$, recompute threshold $\tau_{\rho}$, $k$
@@
4:$\mathcal{S}\leftarrow$ $M$ nearest centroids to $\mathbf{q}$
@@
APS incorporates two optimizations to minimize computational overhead. First, it precomputes values of the regularized incomplete beta function at 1024 evenly spaced points in $[0,1]$ and linearly interpolates during queries. Second, partition probabilities are recomputed only when the query radius $\rho$ shrinks by more than a relative threshold $\tau_{\rho}$. Table 2 shows these optimizations reduce query latency by 25% on SIFT1M without sacrificing recall.
Table 2: Mean single-threaded query latency and recall for APS variants on SIFT1M dataset at recall target 0.9. APS-RP: recomputes probabilities after each partition scan without precomputation. APS-R: recomputes after each partition scan with precomputation. APS: recomputes probabilities only if query radius changes by more than $\tau_{\rho}=1\%$, using precomputed beta function values.
@@ 6 Quake Implementation
To maximize the benefits of data placement, Quake employs partition affinity and NUMA-aware work scheduling. Partitions are bound to specific CPU cores. This binding ensures that partitions are always scanned by the same core to maximize cache utilization. Queries are scheduled to worker threads based on the location of the data partitions they need to access. When a query requires scanning multiple partitions, the work is divided among threads on the relevant NUMA nodes where the partitions reside. By aligning thread execution with data placement, Quake minimizes remote memory accesses and ensures maximum utilization of memory bandwidth.
@@ 7 Experiments
We next evaluate Quake using a number of benchmarks and summarize our main findings:
@@
    Quake achieves the lowest search latency across all dynamic workloads in comparison to state-of-the-art graph indexes, with 1.58×1.5-8$\times$1.5 - 8 × lower search latency than HNSW, DiskANN, and SVS while having 683×6-83$\times$6 - 83 × lower update latency.
@@
    We also find that APS Scanning matches the nprobe of an oracle across recall targets on Sift1M, with only a $17\%$ increase in latency relative to the oracle, and in contrast to existing methods for early termination, APS requires no offline tuning.
@@
    Quake’s NUMA-aware intra-query parallelism exhibits linear scalability and saturates memory bandwidth on the MSTuring100M dataset with 100-million records. Quake achieves 16×16$\times$16 × lower query latency when compared to a single-threaded version and 4×4$\times$4 × lower latency compared to a non-NUMA aware configuration.
@@ 7.1 Workloads
We performed our evaluation on a diverse set of real-world and synthetic workloads, focusing on search and update latency, recall, and scalability.
@@
To test performance under different workload characteristics, we employ a configurable workload generator applicable to any vector dataset. The key parameters include: number of vectors per operation, operation count, operation mix (insert/delete/search ratio), and spatial skew. For skewed workloads, the generator forms clusters and samples queries and updates from these clusters, reflecting hot spots in the vector space.
@@
We also used the vector datasets Sift1M [14] and MSTuring100M [2] to conduct microbenchmarks.
@@ 7.2 Experimental Setup
Large-scale experiments are run on a 4-socket server with Intel Xeon Gold 6148 CPUs (80 cores, 160 threads), 500 GB RAM across 4 NUMA nodes, and  300 GB/s total memory bandwidth. Some microbenchmarks (Tables 2, 5, and LABEL:tab:maintenance_ablation are run on a 2023 MacBook Pro with a M2 Max chip.
Search queries are processed one at a time and we report the total time to process all queries to reach a recall target of 90% recall for $k=100$. Unless otherwise stated, all search numbers use a single worker thread. Quake additionally reports a multi-thread configuration Quake-MT (16 threads) where partition scans are parallelized for individual queries, while Quake-ST uses a single thread for search. For updates we report the total update and maintenance time, where both Quake and the baselines process updates in batches using 16 threads. This setup simulates an online environment where queries arrive individually, and updates are applied in batches. We report maintenance time separately from update latency, as maintenance can be conducted in the background in online systems  [44].
@@
Quake-MT 1.53 .01 .44 1.98 .03 .02 .10 .15 .63 .08 .71 .54 .002 .14 .70
Quake-ST 9.48 .01 .44 9.93 2.43 .08 2.51
Faiss-IVF 165.8 .005 0 $165.8^{+}$45 .01 0 .46 12.25 0 12.25 13.72 0 0 13.74
@@
We configure the main parameters of Quake and the baselines as follows. We disable vector quantization/compression for all baselines, as not all baselines support it. For partitioned indexes we use $sqrt(|X_{0}|)$ partitions where $|X_{0}|$ is the initial number of vectors in the workload. For the graph indexes, we use a graph degree of 64. For LIRE and Quake, we set the partition refinement radius $r=50$. For Quake we set $\tau=250ns$ and use one iteration of k-means for refinement. All systems use 16 threads for updates and maintenance (if applicable). SCANN, DiskANN, and SVS perform maintenance eagerly during an update, therefore we do not measure maintenance time separately from update time. We consider maintenance after each operation for all methods. Throughout all experiments, indexes search parameters are tuned to achieve an average of 90% recall for $k=100$ across the workloads.
@@ 7.3 End-to-End Evaluation
On the OpenImages-13M workload, which includes both insertions and deletions, Quake-MT’s search time is .03 hours. The best competing approach, DiskANN, records .22 hours, making Quake 8×8$\times$8 × faster. Faiss-HNSW does not support deletions so it is omitted. Both SVS’s and DiskANN’s delete consolidation is expensive, leading to multiple orders of magnitude higher update latency than partitioned indexes, illustrating that graph-based indexes struggle with dynamic operations. Quake’s continuous maintenance and adaptive parameter selection keep partitions balanced, achieving low latency and stable recall.
@@
On MSTuring10M-IH, where the dataset grows from one to ten million vectors, Quake-MT achieves a total search time of .54 hours. DiskANN, the second-best performer, has a search time of .81 hours, making Quake 1.5×1.5$\times$1.5 × faster. The other baselines fail to maintain the recall target or suffer from high latency due to their static parameters and inability to prevent partition skew.
Overall, these results demonstrate that Quake’s combination of adaptive partition scanning, incremental maintenance, and NUMA-aware parallelism consistently delivers low-latency queries at the desired recall. Systems without maintenance (Faiss-IVF) suffer from skew-induced latency increases, those tied to static search parameters (LIRE) struggle to maintain recall without incurring higher query times, and graph-based methods (Faiss-HNSW, DiskANN) face substantial overheads when handling updates and deletions. By integrating these components, Quake matches the low update cost of partitioned indexes while outperforming graph indexes in search latency in dynamic workloads. Our design represents a significant advancement to the state-of-the-art, providing stable, efficient performance across diverse, and evolving workloads.
@@
Here we perform a detailed comparison with LIRE and DeDrift, measuring the latency, recall, and number of partitions over time on the Wikipedia-12M workload. For a fair comparison, we disable NUMA to highlight the advantages of APS and maintenance in Quake. The results are shown in Figure 4. First looking at recall, we see that Quake maintains a stable recall of near 90%, while LIRE’s recall degrades over time as it uses a static nprobe. DeDrift’s recall stays relatively constant, as it does not adjust the number of partitions and therefore does not need to adjust nprobe. However, when turning our attention to latency, we see that Quake has near-constant stable latency, even as the dataset grows, while DeDrift’s latency increases significantly with time. In terms of the number of partitions, we see DeDrift stays constant while Quake and Lire increase by 2.5×2.5$\times$2.5 × and 10×10$\times$10 × respectively. LIRE uses significantly more partitions because it uses size thresholding to determine when to split, regardless of whether a given partition is hot or not. Quake on the other hand only splits partitions if their contribution to the cost model is high, allowing for more efficient maintenance. These results show that Quake’s approach to maintenance is superior to existing methods for partitioned index maintenance in minimizing query latency and recall stability.
@@ 7.4 Multi-Query Processing
Here we compare the search performance of Quake and baseline methods in a static batched query setting. Figure 5 shows the measured QPS at a recall of 90%, varying the number of queries in a batch on the Wikipedia-12M workload. The dataset includes all 12M vectors, with 10,000 queries sampled according to Wikipedia page views from December 2021. For Quake, FaissIVF, and SCANN, we use 3,500 partitions; for FaissHNSW, SVS, and DiskANN, we set the graph degree to 64. All experiments use 16 threads for query processing. Quake employs the multi-query execution policy described in [26] and [33], grouping queries by the partitions they access and scanning each partition exactly once per batch in parallel.
Quake consistently outperforms all baselines across every batch size, with an increasing advantage as batch sizes grow. At the largest batch size (10,000 queries), Quake achieves a 6.7x speedup over the partitioned indexes FaissIVF and SCANN. This performance gain stems from Quake’s efficient multi-query execution strategy, where it scans each partition once per batch, in contrast to FaissIVF and SCANN, which scan partitions individually per query. Compared to DiskANN, the strongest graph-based competitor, Quake still maintains a substantial 1.8x speedup. These results demonstrate that Quake delivers high performance not only in single-query scenarios but also in large multi-query workloads.
Configuration Search Latency Recall Std.
Quake-MT .53 ms .008 %
Quake-MT w/o APS .5 ms .025 %
Quake-ST 3.28 ms .005 %
Quake-ST w/o APS 3.18 ms .025 %
Quake-ST w/o Maint/APS 45.2 ms .014 %
@@ 7.5 Scalability
APS 80% 0.821 11.8 0.34ms 0
90% 0.912 20.2 0.48ms 0
99% 0.989 50.1 0.96ms 0
Auncel [47] 80% 0.857 16.4 0.41ms 66.3s
90% 0.981 73.8 1.29ms 73.8s
99% 0.997 95.9 1.61ms 83.2s
SPANN [7] 80% 0.816 11 0.31ms 173s
90% 0.902 19 0.43ms 183s
99% 0.990 70 1.07ms 259s
LAET [18] 80% 0.813 10.5 0.29ms 81s
90% 0.905 18.2 0.42ms 104s
99% 0.990 58.3 1.03ms 232s
Fixed 80% 0.817 11 0.33ms 318s
90% 0.903 19 0.44ms 330s
99% 0.990 65 1.16ms 424s
Oracle 80% 0.833 11.5 0.29ms 320s
90% 0.924 19.3 0.41ms 331s
99% 0.992 42.0 0.74ms 368s

### v2

@@ 4.2.4 Example Maintenance Workflow
    Estimate  For $P_{1}$ and $P_{2}$ the estimate assumes a balanced $/250$ 250 split and $\alpha=0.5$ traffic per child:
@@
    Verify  After splitting we see that $P_{1}$ has a $/250$ 250 split, but $P_{2}$ comes out $450/50$:
@@ 5 Adaptive Partition Scanning (APS)
Adaptive Partition Scanning (APS) dynamically determines the number of partitions to scan per query to achieve a specified recall target $\tau_{R}$ with minimal latency. APS adapts to evolving workloads and changing index structures, making it particularly effective in dynamic data settings. We first introduce the geometric model underlying APS, followed by a detailed description of the scanning algorithm, and conclude with key performance optimizations. We apply APS at each level of the index independently. For clarity we focus on the Euclidean distance, we discuss inner product metrics in the technical report [27].
@@
Exact computation of intersection volumes between a sphere and high-dimensional Voronoi partition boundaries is infeasible, as partitions are intersections of multiple half-spaces. Instead, we approximate each partition as a single half-space, defined by the perpendicular bisector between the query’s nearest centroid $\mathbf{c}_{0}$ and each neighboring centroid $\mathbf{c}_{i}$. This simplification results in a hyperspherical cap whose volume $v_{i}$ has a closed-form expression via the regularized incomplete beta function [19, 16] (see technical report [27]).
@@
$p_{0}=\prod_{j=1}^{M-1}(1-v_{j}),$ (8)
with the remaining probability distributed proportionally among other partitions according to their volumes $v_{i}$:
$p_{i}=(1-p_{0})\,v_{i}.$ (9)
@@ 5.1 APS algorithm
Algorithm 1 details the APS procedure. Given query $\mathbf{q}$, recall target $\tau_{R}$, and the initial candidate fraction $f_{M}$:
@@
This process is conducted at each level of the index. To avoid propagating errors from searching higher levels, we fix the recall target to 99% for the higher levels (Table 6).
@@
1:query $\mathbf{q}$, centroids $\mathcal{C}$, recall target $\tau_{R}$, initial candidate fraction $f_{M}$, recompute threshold $\tau_{\rho}$, $k$
@@
4:$\mathcal{S}\leftarrow$ $f_{M}*N_{l}$ nearest centroids to $\mathbf{q}$
@@
APS incorporates two optimizations to minimize computational overhead. First, it precomputes values of the regularized incomplete beta function at 1024 evenly spaced points in $[0,1]$ and linearly interpolates during queries. Second, partition probabilities are recomputed only when the query radius $\rho$ shrinks by more than a relative threshold $\tau_{\rho}$. Table 2 shows these optimizations reduce query latency by 29% on Sift1M without sacrificing recall.
Table 2: Mean single-threaded query latency and recall for APS variants on Sift1M dataset at recall target 90%. APS-RP: recomputes probabilities after each partition scan without precomputation. APS-R: recomputes after each partition scan with precomputation. APS: recomputes probabilities only if query radius changes by more than $\tau_{\rho}=1\%$, using precomputed beta function values.
@@ 6 Quake Implementation
To maximize the benefits of data placement, Quake employs partition affinity and NUMA-aware work scheduling. Partitions are bound to specific CPU cores. This binding ensures that partitions are always scanned by the same core to maximize cache utilization. Queries are scheduled to worker threads based on the location of the data partitions they need to access. When a query requires scanning multiple partitions, the work is divided among threads on the relevant NUMA nodes where the partitions reside. By aligning thread execution with data placement, Quake minimizes remote memory accesses to maximize memory bandwidth utilization.
@@ 7 Experiments
We evaluate Quake using a number of benchmarks and summarize our main findings:
@@
    Quake achieves the lowest search time across all dynamic workloads compared to state-of-the-art graph indexes, with $1.5$-13×13$\times$13 × lower search latency than HNSW, DiskANN, and SVS while having $18$-126×126$\times$126 × lower update latency.
@@
    We also find that APS matches the nprobe of an oracle across recall targets on Sift1M, with only a $17$-$29\%$ increase in latency relative to the oracle.
@@
    APS performs on-par or better than existing early termination methods [48, 7, 18] and requires no offline tuning.
  4. 4.
    Quake’s NUMA-aware query processing exhibits linear scalability and high memory bandwidth utilization on the MSTuring100M dataset. Quake achieves 20×20$\times$20 × and 4×4$\times$4 × lower query latency compared to single-threaded and non-NUMA aware configurations, respectively.
@@ 7.1 Workloads
We performed our evaluation on a diverse set of real-world and synthetic workloads.
@@
To test performance under varying workload properties, we employ a configurable workload generator applicable to any vector dataset. The key parameters are: number of vectors per operation, operation count, operation mix (read/write ratio), and spatial skew. For skewed workloads, vectors are clustered and sampled from to produce queries and updates, reflecting hot spots in the vector space.
@@
We use the datasets Sift1M [13], Sift10M, and MSTuring100M [2] to conduct microbenchmarks.
@@ 7.2 Experimental Setup
Large-scale experiments are run on a 4-socket server with Intel Xeon Gold 6148 CPUs (80 cores, 160 threads), 500 GB RAM across 4 NUMA nodes, and  300 GB/s total memory bandwidth. Some microbenchmarks (Tables 2, 5, LABEL:tab:maintenance_ablation, and 6) are run on a 2023 MacBook Pro with a M2 Max chip.
Search queries are processed one at a time and we report the total time to process all queries to reach a target of 90% recall for $k=100$. Unless otherwise stated, all search numbers use a single worker thread. Quake additionally reports a multi-thread configuration Quake-MT (16 threads) where partition scans are parallelized for individual queries, while Quake-ST uses a single thread for search. For updates we report the total update and maintenance time, where both Quake and the baselines process updates in batches using 16 threads. This setup simulates an online environment where queries arrive individually, and updates are applied in batches. We report maintenance time separately from update latency, as maintenance can be conducted in the background in online systems  [45].
@@
Quake-MT 1.53 .01 .44 1.98 .03 .02 .10 .15 .63 .08 .71 .54 .02 .14 .70
Quake-ST 9.48 .01 .44 9.93 .14 .02 .10 .26 2.43 .08 2.51 2.12 .02 .14 2.28
Faiss-IVF 165.8 .005 0 $165.8^{+}$45 .01 0 .46 12.25 0 12.25 13.72 .01 0 13.73
@@
We configure the main parameters of Quake and the baselines as follows. We disable vector quantization/compression for all baselines, as not all baselines support it. For partitioned indexes we use $sqrt(|X_{0}|)$ partitions where $|X_{0}|$ is the initial number of vectors in the workload. For the graph indexes, we use a graph degree of 64. For LIRE and Quake, we set the partition refinement radius $r=50$. For Quake we use a single level of partition, set $\tau=250ns$, use one iteration of k-means for refinement and set $f_{M}$ between 1%-10%. All systems use 16 threads for updates and maintenance (if applicable). SCANN, DiskANN, and SVS perform maintenance eagerly during an update, therefore we do not measure maintenance time separately from update time. We consider maintenance after each operation for all methods. Throughout all experiments, indexes search parameters are tuned to achieve an average of 90% recall for $k=100$ across the workloads.
@@ 7.3 End-to-End Evaluation
On the OpenImages-13M workload, which includes both insertions and deletions, Quake’s multi-threaded and single-threaded search times are .03 and .14 hours respectively. The best competing approach, DiskANN, records .22 hours, making Quake-MT 7.3×7.3$\times$7.3 × and Quake-ST 1.6×1.6$\times$1.6 × faster. Faiss-HNSW does not support deletions so it is omitted. Both SVS’s and DiskANN’s delete consolidation is expensive, leading to orders of magnitude higher update latency than partitioned indexes, illustrating that graph-based indexes struggle with dynamic operations. Quake’s continuous maintenance keep partitions balanced, achieving low latency and stable recall.
@@
On MSTuring10M-IH, where the dataset grows from one to ten million vectors, Quake-MT achieves a total search time of .54 hours. DiskANN, the second-best performer, has a search time of .81 hours, making Quake-MT 1.5×1.5$\times$1.5 × faster due to intra-query parallelism. However, single-threaded Quake is 2.6×2.6$\times$2.6 × slower than DiskANN, further illustrating the search efficiency of graph indexes. The other baselines fail to maintain the recall target or suffer from high latency due to their static parameters and inability to prevent partition skew.
Overall, these results demonstrate that Quake’s combination of adaptive partition scanning, incremental maintenance, and NUMA-aware parallelism consistently delivers low-latency queries at the desired recall. Systems without maintenance (Faiss-IVF) suffer from skew-induced latency increases, those tied to static search parameters (LIRE) struggle to maintain recall without incurring higher query times, and graph-based methods (Faiss-HNSW, DiskANN) face substantial overheads when handling updates and deletions. By integrating these components, Quake matches the low update cost of partitioned indexes while outperforming graph indexes in search latency in dynamic workloads. Our design is an advancement to the state-of-the-art, providing stable, efficient performance across diverse, and evolving workloads.
@@
Here we perform a detailed comparison with LIRE and DeDrift, measuring the latency, recall, and number of partitions over time on the Wikipedia-12M workload. For a fair comparison, we use a single-thread to highlight the advantages of APS and maintenance in Quake. The results are shown in Figure 4. First looking at recall, we see that Quake maintains a stable recall of near 90%, while LIRE’s recall degrades over time as it uses a static nprobe. DeDrift’s recall stays relatively constant, as it does not adjust the number of partitions and therefore does not need to adjust nprobe. However, when turning our attention to latency, we see that Quake has near-constant stable latency, even as the dataset grows, while DeDrift’s latency increases significantly with time. In terms of the number of partitions, we see DeDrift stays constant while Quake and Lire increase by 2.5×2.5$\times$2.5 × and 10×10$\times$10 × respectively. LIRE uses significantly more partitions because it uses size thresholding to determine when to split, regardless of whether a given partition is hot or not. Quake on the other hand only splits partitions if their contribution to the cost model is high, allowing for more efficient maintenance. These results show that Quake’s approach to maintenance is superior to existing methods for partitioned index maintenance in minimizing query latency and recall stability.
Configuration Search Latency Recall Std.
Quake-MT 0.53 ms .008
Quake-MT w/o APS 0.50 ms .025
Quake-ST 3.28 ms .005
Quake-ST w/o APS 3.18 ms .025
Quake-ST w/o Maint/APS 45.20 ms .014
@@ 7.4 Multi-Query Processing
Here we compare the search performance of Quake and baseline methods in a static batched query setting. Figure 5 shows the QPS at a recall of 90%, varying the number of queries in a batch on the Wikipedia-12M workload. The dataset includes all 12M vectors, with 10,000 queries sampled according to Wikipedia page views from December 2021. For Quake, FaissIVF, and SCANN, we use 3,500 partitions; for FaissHNSW, SVS, and DiskANN, we set the graph degree to 64. All experiments use 16 threads for query processing. Quake employs the multi-query execution policy in [26] and [34], grouping queries by the partitions they access and scanning each partition exactly once per batch in parallel.
Quake consistently outperforms all baselines across every batch size, with an increasing advantage as batch sizes grow. At the largest batch size (10,000 queries), Quake achieves a 6.7×6.7$\times$6.7 × speedup over FaissIVF and SCANN. This performance gain stems from Quake’s efficient multi-query execution strategy, where it scans each partition once per batch, in contrast to FaissIVF and SCANN, which scan partitions individually per query. Compared to DiskANN, the strongest graph-based competitor, Quake still maintains a substantial 1.8×1.8$\times$1.8 × speedup. These results demonstrate that Quake delivers high performance not only in single-query scenarios but also in large multi-query workloads.
@@ 7.5 Scalability
APS 80% 82.1% 11.8 0.34ms 0
90% 91.2% 20.2 0.48ms 0
99% 98.9% 50.1 0.96ms 0
Auncel [48] 80% 85.7% 16.4 0.41ms 66.3s
90% 98.1% 73.8 1.29ms 73.8s
99% 99.7% 95.9 1.61ms 83.2s
SPANN [7] 80% 81.6% 11 0.31ms 173s
90% 90.2% 19 0.43ms 183s
99% 99.0% 70 1.07ms 259s
LAET [18] 80% 81.3% 10.5 0.29ms 81s
90% 90.5% 18.2 0.42ms 104s
99% 99.0% 58.3 1.03ms 232s
Fixed 80% 81.7% 11 0.33ms 318s
90% 90.3% 19 0.44ms 330s
99% 99.0% 65 1.16ms 424s
Oracle 80% 83.3% 11.5 0.29ms 320s
90% 92.4% 19.3 0.41ms 331s
99% 99.2% 42.0 0.74ms 368s